OpenStudy (chris81):

how can i set this up equal to 0 correctly, logð‘(ð‘£âˆšð‘¢)=logð‘(ð‘¢)/ð‘£

OpenStudy (chris81):

i know its an identity

OpenStudy (chris81):

just rust on how to shift things over

OpenStudy (chris81):

rusty*

OpenStudy (mathmale):

Chris: Could you possibly do a screen shot and share the image with me here?

OpenStudy (chris81):

\[\log_{\sqrt[y]{x}} = \log_{x} /y\]

OpenStudy (mathmale):

Seeing your image made all the difference. So much clearer! You are beginning with \[\log_{b}\sqrt[y]{x} \]

OpenStudy (chris81):

right can i move the stuff over

OpenStudy (chris81):

from the right side

OpenStudy (mathmale):

and want to simplify this expression. First of all, not that this can be re-written as \[\log_{b} x ^{\frac{ 1 }{ y }}\]

OpenStudy (chris81):

to left so it can be \[\log_{b}(\sqrt[y]{x}) (y)(\log(x))\]

OpenStudy (chris81):

=0

OpenStudy (mathmale):

I don't agree that that equals 0. Why are you so sure you're supposed to end up with 0? Please note that your expression, above, is NOT an equation, nor will it become an equation. What do the instructions for this problem say?

OpenStudy (chris81):

oops would it be log(\[\sqrt[y]{x}\])(y)(1/log(x)) = 0

OpenStudy (mathmale):

Hold, please. read the original problem, find the instructions and share them here.

OpenStudy (chris81):

they say they are the same values this values are identities the formula i gave you , so if i move everything over to one side it should give me zero on the other side right?

OpenStudy (mathmale):

You could show that the left side equals the right side. No, that is not right, because this problem does NOT involve addition or subtraction of terms.

OpenStudy (chris81):

oh i see

OpenStudy (chris81):

this makes sence now sorry been really rusty with math lately

OpenStudy (mathmale):

You are beginning with:\[\log_{b}\sqrt[y]{x}\]

OpenStudy (chris81):

right

OpenStudy (mathmale):

As before, you need to re-write that as \[\log_{b}x ^{\frac{ 1 }{ y }} \]

OpenStudy (mathmale):

Do you agree with this? If not, explain why not.

OpenStudy (chris81):

yes i agree

OpenStudy (mathmale):

\[\sqrt[y]{x}=x ^{\frac{ 1 }{ y }}\]

OpenStudy (mathmale):

All right then. If you agree with\[\log_{b}x ^{\frac{ 1 }{ y }}\]

OpenStudy (mathmale):

then we will proceed to apply this rule of logs:\[\log a^b=b*\log a\]

OpenStudy (mathmale):

So: take \[\log_{b}x ^{\frac{ 1 }{ y }}\]

OpenStudy (mathmale):

and rewrite it according to the rule of logs I've just given you.

OpenStudy (chris81):

(1/y)log(x)

OpenStudy (mathmale):

You're not working with log x; you are working with "log to the base b of x. Please fix your response (above).

OpenStudy (mathmale):

You want\[\log_{b}x,~NOT ~\log x \]

OpenStudy (mathmale):

So your \[\frac{ 1 }{ y }\log x\]

OpenStudy (chris81):

\[(1/y)\log_{b}(x) \]

OpenStudy (mathmale):

should be ... \[\frac{ 1 }{ y }\log_{b} x\]

OpenStudy (mathmale):

Yes. That's the right side of your original equation. Work is done. Any questions? Note that there was NO addition or subtraction here.

OpenStudy (chris81):

i see i understand now so i can say if we subtract one side from the other side we should get 0

OpenStudy (mathmale):

No, you can't. Repeat: There is NO addition, NO subtraction, here, only multiplication, division, exponentiation and rules of logs.

OpenStudy (mathmale):

Your job is to show that the equation is true. Forget "addition, subtraction, making it = 0." Not applicable here.

OpenStudy (chris81):

ok

OpenStudy (mathmale):

Hope you are satisfied. Post another question (separately), if you like.